# Emergent active turbulence and intermittency in dense algal suspensions of Chlamydomonas reinhardtii

### v2

@@ II Statistical characterisation of dense algal suspensions
We turn now to an investigation of possible signatures of an active glass in our WT and mbo2 systems at large values of $\bar{\rho}$. Figures 5 (a), (b), and (c) show, respectively, plots of the mean-square displacement (MSD), the modulus $|F_{s}(k_{x},t)|$ of the self-intermediate-scattering function (for $k_{x}a_{w}=8.35$), the self-overlap function $Q(t)$, and the four-point correlation function $\chi_{4}(t)$ versus the scaled time $\Delta t/\tau_{w}$ [see Eqs.(12), (13), (14), (15)] for WT C. reinhardtii cells with $\bar{\rho}=0.74$; the reference time scale $\tau_{w}\equiv a_{w}/v_{w}$, where $v_{w}\simeq 100\;\mu{\rm{m/s}}$ is the mean swimming speed of isolated WT cells; the local slope $\frac{\delta(log\langle\Delta x^{2}\rangle)}{\delta(log\Delta t)}$ is shown in the inset of Fig. 5 (a). Figures 5 (d), (e), and (f) are the mbo2 counterparts of Figs. 5 (a), (b), and (c); for mbo2, $k_{x}a_{m}=15.41$ and the reference time scale $\tau_{m}\equiv a_{m}/v_{m}$, where $v_{m}\simeq 50\;\mu{\rm{m/s}}$ is the mean swimming speed of isolated mbo2 cells. We observe that there are no well-developed plateaux in the plots of the MSD, $F_{s}(k,t)$, and $Q(t)$; such plateaux are desiderata for an active glass  [23]. However, the presence of a peak in the four-point correlation function $\chi_{4}(t)$ suggests dynamic heterogeneity in our algal systems.
III Algal turbulence versus an active algal glass
We have uncovered a new type of emergent active turbulence in dense algal systems of two types of C. reinhardtii cells. This algal turbulence displays power-law energy spectra and signatures of small-scale intermittency that are qualitatively similar to their counterparts in conventional fluid turbulence [see, e.g., Refs. [5, 4]] and in bacterial turbulence, which has garnered considerable attention over the past decade [see, e.g., Refs. [9, 7, 27, 13]]. However, there are important quantitative differences between the algal turbulence, which we investigate, and fluid and bacterial turbulence. We have characterised these differences by computing various statistical properties of algal turbulence.
Microswimmer turbulence is often associated with the presence of topological defects and orientational instabilities and, therefore, is limited to rod-shaped bacteria. Our study shows that turbulence can also arise in the absence of orientational instabilities and hence is a universal feature of free-swimming micro-organisms, regardless of their shape.
Reference [26], which has carried out a theoretical and numerical study of the active Cahn Hilliard Navier Stokes (CHNS) equations, suggests that active-scalar turbulence in this system might be realised in dense suspensions of C. reinhardtii. The two-dimensional (2D) active CHNS model they consider for a dense suspension of contractile swimmers [32, 15] is:
$\partial_{t}\phi+(\bm{u}\cdot\nabla)\phi$ $=$ $\text{M}\nabla^{2}\left(\frac{\delta\mathcal{F}}{\delta\phi}\right)\,;$ (1)
$\partial_{t}\omega+(\bm{u}\cdot\nabla)\omega$ $=$ $\nu\nabla^{2}\omega+\frac{3}{2}\epsilon\nabla\times(\nabla\cdot\bm{\Sigma}^{A})-\alpha\omega\,;\;$ (2)
$\nabla\cdot\bm{u}$ $=$ $0\,;$ (3)
the vorticity $\omega=\nabla\times\bm{u}$ and M, $\nu$, and $\alpha$ are, respectively, the mobility, kinematic viscosity, and bottom friction, and the Landau-Ginzburg variational free-energy functional is
$\mathcal{F}[\phi,\nabla\phi]=\int_{\Omega}\left[\frac{3}{16}\frac{\sigma}{\epsilon}(\phi^{2}-1)^{2}+\frac{3}{4}\sigma\epsilon|\nabla\phi|^{2}\right]\,;$ (4)
the first term in the integrand has minima at $\phi=\pm 1$; and the scalar $\phi$ is, respectively, positive where the microswimmer density is high and negative where this density is low; $\phi$ varies smoothly across interfaces between these regions. The bare surface tension $\sigma$ measures the free-energy cost for an interface, whose width $\sim\epsilon$. In this active model, all terms in the stress tensor cannot be derived from $\mathcal{F}$; the stress tensor $\bm{\Sigma}^{A}$ has the components [15, 32, 33, 34]
$\Sigma^{A}_{ij}=-\zeta\left[\partial_{i}\phi\partial_{j}\phi-\frac{\delta_{ij}}{2}|\nabla\phi|^{2}\right]\,;$ (5)
$\zeta$, the activity coefficient, is negative for contractile swimmers [32] but positive for their extensile counterparts. The density in the algal suspension depends on $\phi$; in particular, $\rho(\phi)\equiv\rho_{1}(1+\phi)/2+\rho_{2}(1-\phi)/2$, where $\rho_{1}$ and $\rho_{2}$ are the densities of the algal-rich and algal-poor regions.
A comparison of our results for $E(k)$ and $\Phi(k)$ with those of Ref. [26] [see Figs. 3(a) and (b) there] shows a qualitative similarity between these spectra, but the exponents, which characterise the power-law spectral regimes, are not the same. Furthermore, the active-CHNS system yields a turbulent state only for contractile swimmers; here, we find turbulence for both the WT and mbo2 cells; this discrepancy might arise because WT and mbo2 cells are contractile or extensile only on average, as the flagellar stroke during a beat cycle varies [35]. Therefore, the theoretical modelling of algal turbulence in dense suspensions of C. reinhardtii cells is still a challenging, open problem; to the extent that our results for WT and mbo2 variants are qualitatively similar, we might well have to go beyond models, like the one proposed in Ref. [26], that yield qualitatively different results for contractile and extensile microswimmers.
@@ III Algal turbulence versus an active algal glass
In summary, we have demonstrated that, at low values of $\bar{\rho}$, our algal systems display an emergent nonequilibrium state with statistically steady algal turbulence. Although this algal turbulence decreases, with increasing $\bar{\rho}$, it does not cross over completely to a state that can be identfied as a bona fide active algal glass. It might well turn out that our dense suspensions of C. reinhardtii cross over to an active algal glass at values of $\bar{\rho}$ that are larger than those we have considered here; however, such high densities are hard to achieve in experiments, so the explorations of active glassy dynamics in very dense suspensions of C. reinhardtii remains a challenge for future studies. Furthermore, algal turbulence has obvious biological implications; such turbulence enhances mixing and transport, so we conjecture that it aids C. reinhardtii in foraging for and outrunning diffusing nutrient molecules.
@@ IV.1 Cell culture
Chlamydomonas reinhardtii cells of both the WT and mbo2 strains are inoculated from agar plates into TAP+P media and grown at 25°$\degree$ Celsius in 12:12 hour day-night cycles inside an orbital shaker at 137rpm [36]. When the cells are in the logarithmic phase of their growth cycle, the culture is collected in Eppendorf tubes and then centrifuged multiple times (at 0.3rcf for 4 minutes) to obtain samples with different cell concentrations.
C. reinhardtii is a photosynthetic alga that feeds on dissolved inorganic ions or molecules that are in the TAP+P media; these include phosphate and ammonium ions and carbon dioxide from the surrounding fluid; light is the main source of energy [see, e.g., Refs.  [37, 38]]. The macronutrients that limit algal metabolism and growth are nitrogen and carbon [see, e.g., Refs.  [39, 40, 38]]. Flow fields that are flagella-driven help to distribute these dissolved solute molecules uniformly through fluid mixing and transport [see, e.g., Refs.  [38, 37, 41, 40, 17, 19]].
IV.2 Surface modification of glass slides, cover slips and beads
A polyacrylamide brush is coated on the glass slides and cover slips to avoid non-specific adhesion of cells and beads on the glass surfaces [42]. Also, the microspheres (Sulphate latex 200nm beads, Thermofisher) are coated with PLL-PEG to impart steric stabilization, thereby reducing inter-particle aggregation and obtaining mono-dispersed microspheres in the media [42].
@@ IV.3 Microscopic imaging
Double-sided tape of height 10 micron (Nitto Denko Corporation) is used as a spacer between the coverslip and the glass slide to obtain the chambers in which the cell-culture is placed (Supplementary Information Section 2). The sample is kept under an inverted microscope (Olympus IX83) coupled with a CMOS high-speed camera (Phantom Miro C110, Vision Research, Pixel size = 5.6 micron). Different concentrations of both the WT and mbo2 mutant cells are imaged under red-light ($>610$nm) using a 20x objective in bright field at 100 fps. Dilute cultures are mixed with 200nm microspheres and then imaged using a 60x phase objective at 500 fps to obtain flow fields of isolated cells  [43].
IV.4 Obtaining PIV, Cell Density, and Flow-fields from the images
The velocity vectors are obtained using the Matlab tool PIVlab [44] with a time interval of 50ms between successive frames. We use the FFT window deformation algorithm and two passes to analyze the images. For the first pass, we use an interrogation window of size $16.8\mu m$, and the second pass has a size of $8.4\mu m$. We follow the bacterial-suspension studies of Refs. [9, 45], in which bacteria are used as tracers; here, PIV tracks bacterial flow, instead of the fluid flow; nevertheless, the fluid and bacterial flows have been found to share quantitatively similar statistical structures [9, 45]. We use PIV to track algal flows and assume, as in Refs. [9, 45], that the algal and fluid flows are statistically similar. The spatial cell density is determined by calculating the area fraction covered by the cells at a particular location for each frame (Fig.3, Supplementary Information). Particle tracking velocimetry (PTV), wherein individual microspheres are tracked to calculate their velocity vectors, was used to compute the beat averaged flow fields of isolated WT and mbo2 cells  [43].
@@ IV.5 Statistical measures for algal turbulence
    here, $\langle\cdot\rangle_{t}$ is the time average; typically we take the mean over $\simeq 800$ spectra, obtained from configurations of the density and velocity fields, at well-separated times in the nonequilibrium statistically steady state. We also average over different samples with similar densities [see Fig. 2].
@@
    for different length scales $l$.
  • Furthermore, we compute the Okubo-Weiss parameter [28, 29, 5]
@@
    $D$ and $\Omega$ are the rate-of-deformation and rate-of-rotation tensors, respectively; $\Lambda>0$ in strain-dominated regions of the flow and $\Lambda<0$ in vortical regions.
  • We compute the flatness $F_{4}(l)$ to check for intermittency in our system  [1, 46, 47, 5]:
@@ IV.6 Statistical measures for active glasses
In the active-glass literature [see, e.g., Ref. [23]] various statistical properties are used to characterise a glass. The analysis that we present here is based on the particles (here, C. reinhardtii cells) that comprise the glass; by contrast, our analysis of turbulence in the previous Section is based on density and velocity fields of the type that are used in hydrodynamical models of active-scalar turbulence [26]. The statistical properties that we use to explore the formation of an active glass in our system are given below:
@@
    here, $\bm{x}_{i}(t)$ is the position of the $i^{th}$ cell at time $t$ and $N$ is the total number of cells. We use $N\simeq 1000-1500$.
@@
V Acknowledgments
P.V.B and P.S thank the ANRF grant CRG/2022/000724 and the CSIR grant 03WS(008)/2023-24/EMR-II/ASPIRE for providing the support to carry out the experiments. N.B.P., B.M,, and R.P thank K.V. Kiran for discussions, the Anusandhan National Research Foundation (ANRF), the Science and Engineering Research Board (SERB), and the National Supercomputing Mission (NSM), India, for support, and the Supercomputer Education and Research Centre (IISc), for computational resources.
Supplementary Information
This Supplementary Information contains some details about the cells sizes of the two strains, height measurement of the sample chamber, fitted PDFs of the $x$ component of the velocity and the longitudinal velocity increments, computations of the spatial density from the microscopic image, and a list of videos that show the spatiotemporal evolution of the CR cells and also of the pseudocolor plots in Figs. 1 (g), (h), (j), and (k) [in the main paper].
I Cell sizes of different strains
Figure 1 presents histograms of the mean diameters for both the Wild-type (WT) and the mbo2 mutant C. reinhardtii cells. The average diameter of the WT cell is $a_{w}=7.82\pm 0.84\ \mu$m and for the mbo2 cell $a_{m}=7.22\pm 0.85\,\mu$m.
@@ I Cell sizes of different strains
II Height measurement of the chamber
We take a dilute solution of microspheres (Sulphate latex 200nm beads, from Thermofisher) on the coverslip and the glass slide; we then heat the glass slide and the coverslip gently so that the beads get stuck on both the surfaces. After that, immersion oil is put into the sample chamber. We use a 60x oil-immersion phase objective to focus on the beads and determine the chamber height to be $10.27\pm 0.61\ \mu$m.
III Fitted PDFs
Figure 2: Fitted PDFs of (a) the $x$-component of the velocity and (b) the longitudinal velocity increments for the WT cells. (c) and (d) are the mbo2 counterparts of (a) and (b).

### v3

@@ II Statistical characterisation of dense algal suspensions
Figure 6: Plots versus the scaled time $\Delta t/\tau_{w}$ of (a) the mean-square displacement (MSD) [the local slope $\frac{\delta(log\langle\Delta x^{2}\rangle)}{\delta(log\Delta t)}$ is shown in the inset], (b) the modulus $|F_{s}(k_{x},t)|$ of the self-intermediate-scattering function (for $k_{x}a_{w}=8.38$), and (c) the self-overlap function $Q(t)$ and the four-point correlation function $\chi_{4}(t)$ [Eqs.(7), (8), (9), (10)] for WT C. reinhardtii cells with $\bar{\rho}=0.75\pm 0.02$; the reference time scale $\tau_{w}$ is defined in the text; (d), (e) and (f) are the mbo2-mutant counterparts of those in (a), (b) and (c), respectively. Our data are averaged over cell-trajectories of samples with similar values of $\bar{\rho}$ (within $\simeq 10\%$ of the mean); the error bars denote $\pm\varsigma(\Delta x^{2}(t))$ and $\pm|\varsigma(F_{s}(k_{x},t))|$.
Note that the PDFs of $\bm{u}_{x}$ are markedly different from those that are obtained for conventional fluid turbulence pandit2017overview and bacterial turbulence in dense suspensions of Bacillus subtilis dunkel2013fluid, both of which are Gaussian. By contrast, the PDFs of $\bm{u}_{x}$, in Figs. 4(a) and (c), are distinctly non-Gaussian (we fit them to compressed exponentials in Appendix I). The non-normalized flow speeds (Appendix F) decrease with an increase in cell concentration. Furthermore, $\mathcal{P}(\Lambda)$ in the algal-turbulence state of our dense suspensions of C. reinhardtii cells is significantly different from its counterparts in 2D fluid turbulence Perlekar2009-kk and bacterial turbulence KiranPRF2023.
Figure 5(a) [blue panel] displays the PDF ($\mathcal{P}$) of the longitudinal velocity increments $\delta\bm{u}_{\parallel}$(l) of WT cell suspensions for different mean densities $\bar{\rho}$ and the separation $l/a_{w}=5.37$ (Materials and Methods, Eq. (4pandit2017overview). Figure 5(c) [beige panel] gives the corresponding PDF for the mbo2 mutant (here $l/a_{m}=5.82$). The $l$-dependence of these velocity-increment PDFs are shown in Appendix H. This $l$-dependence is a fingerprint of intermittency, which we quantify by the flatness $F_{4}$ that is related to the fourth- and second-order velocity structure functions [Materials and Methods, Eq. (6)]. We plot the flatness $F_{4}$, for both WT (versus $l/a_{w}$) and mbo2 (versus $l/a_{m}$) variants, in Figs. 5 (b) and 5 (d), respectively. Both these plots show distinct deviations from the Gaussian value $F_{4}^{G}=3$; this deviation increases as $l/a_{w}$ (or $l/a_{m}$) decreases, a clear manifestation of small-scale intermittency in the algal turbulence we consider  [^1] Signatures of small-scale intermittency are often uncovered by such plots of flatness or the hyperflatness $F_{6}\equiv S_{6}/(S_{2})^{3}$ in fluid and other forms of turbulence Frisch-CUP; perlekar2006manifestations; sahoo2011systematics; pandit2017overview.
We turn now to an investigation of possible signatures of an active glass in our WT and mbo2 systems at large values of $\bar{\rho}$. Figures 6 (a), (b), and (c) show, respectively, plots of the mean-square displacement (MSD), the modulus $|F_{s}(k_{x},t)|$ of the self-intermediate-scattering function (for $k_{x}a_{w}=8.38$), the self-overlap function $Q(t)$, and the four-point correlation function $\chi_{4}(t)$ versus the scaled time $\Delta t/\tau_{w}$ [Eqs. (7), (8), (9), and (10)] for WT C. reinhardtii cells with $\bar{\rho}=0.75$; the reference time scale $\tau_{w}\equiv a_{w}/v_{w}$, where $v_{w}\simeq 100\;\mu{\rm{m/s}}$ is the mean swimming speed of isolated WT cells; the local slope $\frac{\delta(log\langle\Delta x^{2}\rangle)}{\delta(log\Delta t)}$ is shown in the inset of Fig. 6 (a). Figures 6 (d), (e), and (f) are the mbo2 counterparts of Figs. 6 (a), (b), and (c); for mbo2, $k_{x}a_{m}=15.47$ and the reference time scale $\tau_{m}\equiv a_{m}/v_{m}$, where $v_{m}\simeq 50\;\mu{\rm{m/s}}$ is the mean swimming speed of isolated mbo2 cells. We observe that there are no well-developed plateaux in the plots of the MSD, $F_{s}(k,t)$, and $Q(t)$; such plateaux are desiderata for an active glass  sadhukhan2024perspective. However, the presence of a peak in the four-point correlation function $\chi_{4}(t)$ suggests dynamic heterogeneity in our algal systems. In Appendix G, we show the variation, with respect to the average cell-density, of the intermediate scattering function, the overlap function, the 4-point correlation function and the MSD; and how the intermediate scattering function varies for different values of $k$. We note that the MSD plots show diffusive behaviors at short time scales, which is typical of crowding in dense suspensions Koorehdavoudi2017-lz; at longer time scales, the MSD show super-diffusive behavior, as expected Ariel2015-dc; mukherjee2021anomalous.
III Discussion
@@
In summary, we have demonstrated that dense suspensions of Chlamydomonas reinhardtii algal cells display a new type of emergent active turbulence with strongly non-Gaussian velocity distributions, a feature distinct from both bacterial and classical fluid turbulence  pandit2017overview; boffetta2012two; dunkel2013fluid; alert2020universal; mukherjee2023intermittency; KiranPRF2023 . This algal turbulence also displays power-law energy spectra and small-scale intermittency that are quantitatively different from their counterparts in conventional fluid turbulence and in bacterial turbulence . Although the overall dynamics in algal suspensions slows down with increasing number density of cells, we do not observe a cross over to a state that can be identified as a bona fide active algal glass. It might well turn out that an active algal glass appears at densities values larger than those we have considered here; however, such high densities are hard to achieve in experiments, so the explorations of active glassy dynamics in very dense suspensions of C. reinhardtii remains a challenge for future studies.
Furthermore, algal turbulence has obvious biological implications; such turbulence enhances mixing and transport, so we conjecture that it aids C. reinhardtii in foraging for and outrunning diffusing nutrient molecules debasmita_flowinversion. Microbial turbulence is often associated with the presence of topological defects and orientational instabilities. Our study shows that turbulence can also arise in the absence of orientational instabilities. Hence, we hypothesize that active turbulence is a universal feature of free-swimming micro-organisms, regardless of their shape.
Current theoretical models of active turbulence are either inappropriate or insufficient to explain our results. The Toner-Tu-Swift-Hohenberg (TTSH) model bratanov2015new; wensink2012meso, which is commonly used for bacterial turbulence, cannot be applied to algal system because the C. reinhardtii cells do not swarm or flock; and they have an aspect ratio very close to one (unlike rod-like bacteria). Recent work on theoretical and numerical study of the active Cahn Hilliard Navier Stokes (CHNS) equations suggests that active turbulence might be realized in dense suspensions of active-scalar matter  padhan2024novel. However, this active-CHNS model also does not seem to be adequate for the description of our algal system, in so far as the statistical properties of the algal turbulence are different from those in the model. Therefore, the theoretical modeling of algal turbulence in dense suspensions of C. reinhardtii cells remains a challenging open problem.
@@ IV.1 Cell culture
[…]
C. reinhardtii is a photosynthetic alga that feeds on dissolved inorganic ions or molecules that are in the TAP+P media; these include phosphate and ammonium ions and carbon dioxide from the surrounding fluid; light is the main source of energy Tam2011-rz; Kiørboe2008amechanical. The macronutrients that limit algal metabolism and growth are nitrogen and carbon Khan2018-vm; Short2006-dj; Kiørboe2008amechanical. Flow fields that are flagella-driven help to distribute these dissolved solute molecules uniformly through fluid mixing and transport Kiørboe2008amechanical; Tam2011-rz; Ding2014-xx; Short2006-dj; leptos2009dynamics; gollub2011enhancement.
IV.2 Surface modification of glass slides, cover slips, and beads
A polyacrylamide brush is coated on the glass slides and cover slips to avoid non-specific adhesion of cells and beads on the glass surfaces debasmita_flagella. Also, the microspheres (Sulphate latex 200nm beads, Thermofisher) are coated with PLL-PEG to impart steric stabilization, thereby reducing inter-particle aggregation and obtaining mono-dispersed microspheres in the media debasmita_flagella.
Despite the surface functionalization, a small fraction of cells get stuck on the glass surface. Samples of the same cell-type and concentration typically have a percentage of stuck cells that varies between $1\%$ and $4\%$. This small proportion of stuck cells does not play a significant role in the measured statistics of the active turbulence in the system.
@@ IV.3 Microscopic imaging
Double-sided tape of height 10 micron (Nitto Denko Corporation) is used as a spacer between the coverslip and the glass slide to obtain the chambers in which the cell-culture is placed (Appendix B). The sample is kept under an inverted microscope (Olympus IX83) coupled with a CMOS high-speed camera (Phantom Miro C110, Vision Research, Pixel size = 5.6 micron). Different concentrations of both the WT and mbo2 mutant cells are imaged under red-light (>610>610nm) using a 20x objective in bright field at 100 fps. Dilute cultures are mixed with 200nm microspheres and then imaged using a 60x phase objective at 500 fps to obtain flow fields of isolated cells  debasmita_flowinversion.
IV.4 Obtaining Velocity vectors, Cell Density, and Flow-fields from the images
The velocity vectors are obtained by tracking the individual cells between successive frames using standard MATLAB tracking routines  blair2008matlab. For dense suspensions, frames are skipped so that there is sufficient displacement between successive frames and the tracking is accurate. To compute the energy spectra and the Okubo-Weiss parameter, the velocity vectors (obtained from discrete individual cells) are interpolated over a finer equally-spaced grid via Natural-neighbor interpolation. For all other quantities that we compute, we use the non-interpolated velocity vectors.
The spatial cell density is determined by calculating the area fraction covered by the cells at a particular location for each frame (Appendix J). To compute the beat-averaged flow fields of isolated WT and mbo2 cells, the individual microspheres are tracked and their velocity vectors are computed debasmita_flowinversion.
@@ IV.5 Statistical measures for algal turbulence
    here, $\langle\cdot\rangle_{t}$ is the time average; typically we take the mean over $\simeq 100$ energy spectra and $\simeq 400$ concentration spectra, obtained from configurations of the velocity and density fields, at well-separated times in the nonequilibrium statistically steady state. We also average over different samples with similar densities [Fig. 3].
@@
  • Furthermore, we compute the Okubo-Weiss parameter weiss1991dynamics; okubo1970horizontal; pandit2017overview
@@
    $D$ and $\Omega$ are the rate-of-deformation and rate-of-rotation tensors, respectively; $\Lambda>0$ in vortical regions of the flow and $\Lambda<0$ in strain-dominated regions.
  • We compute the flatness $F_{4}(l)$ to check for intermittency in our system  Frisch-CUP; perlekar2006manifestations; sahoo2011systematics; pandit2017overview:
@@ IV.6 Statistical measures for active glasses
In the active-glass literature sadhukhan2024perspective, various statistical properties are used to characterise a glass. The analysis that we present here is based on the particles (here, C. reinhardtii cells) that comprise the glass. The statistical properties that we use to explore the formation of an active glass in our system are given below:
@@
    here, $\bm{x}_{i}(t)$ is the position of the $i^{th}$ cell at time $t$ and $N$ is the total number of cells. We use $N\simeq 1000$.
@@
Acknowledgements.
P.V.B and P.S thank Debasmita Mondal for sharing the PTV codes; Tejas G Murthy, Aparna Baskaran, Sriram Ramaswamy and Chandan Dasgupta for their valuable discussions; the ANRF grant CRG/2022/000724 and the CSIR grant 03WS(008)/2023-24/EMR-II/ASPIRE for providing the support to carry out the experiments. N.B.P, B.M, and R.P thank K.V. Kiran for discussions, the Anusandhan National Research Foundation (ANRF), the Science and Engineering Research Board (SERB), and the National Supercomputing Mission (NSM), India, for support, and the Supercomputer Education and Research Centre (IISc), for computational resources.
@@ Acknowledgements.
Appendix A Cell sizes of different strains
[…]
Appendix B Height measurement of the chamber
[…]
Appendix C Orientational order
The orientation angle, $\theta$ of a CR cell is the angle of the instantaneous velocity vector of a cell relative to the $x$ axis in the lab frame. Figure 8(a) shows a schematic of the motion of a cell; the orientation angle, $\theta$ and the velocity vector, $v$ are shown along its trajectory. Figure 8(b) shows the average polarization $P(t)=<\cos\theta>$ versus time t.
Figure 8: (a) Orientation angle, $\theta$ of a CR cell. (b) Average polarization of the cell suspension as a function of time t, for wild-type (WT) cells and representative $\bar{\rho}=0.50$.
Figure 9: (a) The orientation vectors of the CR cells, over a microscope image with a superimposed spatial grid; (b) and (c) show, respectively, the probability distribution functions (PDFs) $f(r)$ of the orientational order parameter $|<e^{i\theta}>_{grid}|$ and the nematic order parameter $|<e^{2i\theta}>_{grid}|$ averaged over different grid sizes.
Figure 10: The probability distribution functions (PDFs) $f(r)$ of (a) the orientational order parameter $|<e^{i\theta}>_{grid}|$ and (b) the nematic order parameter $|<e^{2i\theta}>_{grid}|$ for WT cells and $\bar{\rho}=0.74$
Figure 9(a) shows the orientation vectors of the CR cells ($\bar{\rho}=0.50$), over a representative image of the cell suspension with a superimposed spatial grid. Figures 9(b) and (c) present, respectively, the probability distribution functions (PDFs) of the orientational order parameter $|<e^{i\theta}>_{grid}|$ and its nematic counterpart $|<e^{2i\theta}>_{grid}|$ [Eq. (11)]. The subscript gridgrid indicates an average over cells within our grids and we consider several grid sizes. The probability distribution function (PDF) $f(r)$ is defined as
$f(|<e^{ki\theta}>_{grid}|=r)=\frac{1}{2\pi r}\int_{0}^{2\pi}pdf(|<e^{ki\theta}>_{grid}|=r,arg(<e^{ki\theta}>_{grid})=\phi)d\phi\,.$ (11)
Had there been any peaks in the PDFs in Figs. 9(b) and (c) they would have signified orientational or nematic order. Similar plots of the probability distribution function (PDF) $f(r)$ for $\bar{\rho}=0.74$ is shown in Fig. 10.
Appendix D Spatial velocity-velocity correlation
Figure 11 shows the spatial velocity-velocity correlation function, $\frac{<v(x).v(x+\Delta l)>}{<v(x)^{2}>}$ for WT cells for two different densities. We fit the initial part of the graph with an exponential function, $\frac{<v(x).v(x+\Delta l)>}{<v(x)^{2}>}=e^{-(\Delta l/a_{w})/\lambda}$ to estimate the correlation length $\lambda$.
Figure 11: Spatial velocity-velocity correlation function for WT cells versus the scaled distance $\Delta l/a_{w}$ at two representative values of $\bar{\rho}$ (a) 0.50 and (b) 0.74. Insets show exponential-decay fits to the initial part of the correlation functions. The decay correlation length $\lambda=0.54$ for $\bar{\rho}=0.50$, and $\lambda=0.63$ for $\bar{\rho}=0.74$, indicating that velocity correlations die out very rapidly (even below one cell-length).
Appendix E Characteristic length scales
We plot two natural characteristic length scales $L_{E}$ and $L_{\phi}$ as a function of time in Fig. 12. The length scales for different cell-densities are shown in Fig. 13.
Figure 12: The temporal evolution of the characteristic length scales (a) $L_{E}$ and (b) $L_{\phi}$ for WT cells (for $\bar{\rho}=0.50$).
Figure 13: The characteristic length scales (a) $L_{E}$ and (b) $L_{\phi}$ for WT cells of different cell-densities. Panels (c) and (d) are the mbo2 counterparts of (a) and (b), respectively.
Figure 14: Semi-log plots of the PDFs of the $x$-component of velocity for (a) WT cells (blue background) and (b) mbo2 cells (beige background) showing the flow speeds in $\mu m/sec$.
Figure 15: The intermediate scattering function (top panel), the overlap function along with the 4-point correction function (bottom panel) is shown for three different cell densities.
Figure 16: (a) The area under the scattering function and (b) the peak in the 4-point correlation function plotted for three different cell-densities. The dashed line is a guide to the eye.
Figure 17: Log-log plots of the mean-square displacement (MSD) for different cell concentrations; the error bars denote $\pm\varsigma(\Delta x^{2}(t))$.
Figure 18: Plots of the intermediate scattering function for different values of $k_{x}a_{w}$ (WT cells with $\bar{\rho}=0.75\pm 0.02$).
Appendix F PDFs of velocity components
We show the PDFs of the $x$ component of the velocity with the non-normalized flow speeds in Fig. 14.
Appendix G Active-glass characterization
The evolution of the intermediate scattering function, the overlap function and the 4-point correlation function with cell-density is shown in Fig. 15. Figure 16(a) shows that the area under the curve of the intermediate scattering function increases with the average cell-density. Similarly, the peak position of the 4-point correlation function in Fig. 16(b) increases with an increase in the density. Figure 17 shows the mean-square displacement (MSD) for different concentrations of WT cells.
The intermediate scattering function for different values of $k$ are shown in Figure. 18. As $k$ increases, there is a shift in the function. However, the overall behavior of the function remains the same.
Appendix H PDFs of velocity increments
Figure 19 displays PDFs ($\mathcal{P}$) of the longitudinal velocity increments $\delta\bm{u}_{\parallel}$(l) of WT cell suspensions for different values of $l$ and $\bar{\rho}$. Figure 20 gives the corresponding PDFs for the mbo2 mutant.
Figure 19: Semilog plots, for WT cells, of the PDFs of longitudinal-velocity increment of the velocity field $\bm{u}$, for different values of separation $l$ and mean density $\bar{\rho}$.
Figure 20: Semilog plots, for mbo2 cells, of the PDFs of longitudinal-velocity increment of the velocity field $\bm{u}$, for different values of separation $l$ and mean density $\bar{\rho}$.
Appendix I Fitted PDFs